OpenStudy (anonymous):

One of the factors of g2 - 32g + 256 is g + 8 g - 8 g - 16 g + 16

OpenStudy (anonymous):

@triciaal is it g-16

OpenStudy (danjs):

factor g^2 - 32g + 256

OpenStudy (triciaal):

yes

OpenStudy (danjs):

(g+8)(a*g+b)

OpenStudy (triciaal):

(a-b)^2= a^2 -2ab + b^2

OpenStudy (michele_laino):

please note that: \[g ^{2}-32g+256=(g-16)^{2}\]

OpenStudy (michele_laino):

it is also: \[g ^{2}-32g+256=(g+8)(g-40)+576\] if we perform the division between polynomials